# Decoherence of Histories: Chaotic Versus Integrable Systems

### v1

@@ Abstract
Introduction.—The decoherence functional (DF)—quantifying the (de)coherence between different paths or histories in an isolated quantum system (precisely defined below)—plays a crucial role to explain the emergence of classicality [1, 2, 3, 4]. Historically, owing to the complexity of the DF, research has been restricted to evaluating it explicitly for the case of quantum Brownian motion only [5, 6, 7, 8, 9, 10, 11]. This non-interacting integrable model has been solved for a bath prepared in a canonical ensemble, a state that is highly mixed and contains (in the conventionally considered continuum limit) an infinite amount of classical noise. This leaves it unclear whether the microscopic origin of the observed decoherence is an intrinsic feature of the dynamics or an artifact of the classical initial state. Also other indirect arguments for emergent decoherent histories were based on linear oscillator chains [12]. In addition, the thermodynamic limit makes the question how exactly decoherence emerges as a function of system size inaccessible.
@@
Our results further shed light on the debated relation between environmentally induced decoherence (EID) [25, 26, 27] and decoherent histories [28, 29, 30, 31, 32, 21]. In our model, the relevant reduced density matrix exactly commutes with the relevant observable at all times, yet the histories are not exactly decoherent as one might naively expect. Since the block-diagonal form of the reduced density matrix is in our example caused by symmetry, this does not contradict the idea of EID, but it illustrates how subtle the relation is: it is a clear-cut example for emergent decoherent histories that are not caused by the entanglement between subsystems.
@@
Preliminaries.—We consider an isolated quantum system with Hamiltonian $H$, Hilbert space ${\cal H}$ with dimension $D=\dim{\cal{H}}$ and initial state $|\psi_{0}\rangle$. We divide the Hilbert space ${\cal{H}}=\bigoplus_{x}{\cal{H}}_{x}$ into orthogonal subspaces ${\cal{H}}_{x}$ corresponding to a complete set of orthogonal projectors $\{\Pi_{x}\}_{x=1}^{M}$ satisfying $\Pi_{x}\Pi_{y}=\delta_{x,y}\Pi_{x}$ and $\sum_{x=1}^{M}\Pi_{x}=I$ (with $I$ the identity). We call the set $\{\Pi_{x}\}$ a coarse-graining because in applications related to the emergence of classicality the projectors $\Pi_{x}$ belong to some coarse observable $A=\sum_{x=1}^{M}a_{x}\Pi_{x}$ characterized by large subspace dimensions $V_{x}\equiv\dim{\cal{H}}_{x}\gg 1$.
@@
Here, $U_{k,i}=e^{-iH(t_{k}-t_{i})}$ is the unitary time evolution operator from time $t_{i}$ to $t_{k}$ ($\hbar\equiv 1$). For brevity, we denote a history as $\boldsymbol{x}=(x_{n},\dots,x_{1},x_{0})$ such that Eq. (1) becomes $|\psi_{n}\rangle=\sum_{\boldsymbol{x}}|\psi(\boldsymbol{x})\rangle$. Moreover, the length of a history is given by $L=n+1$.
@@
which quantifies the overlap, or interference, between different histories $\boldsymbol{x}$ and $\boldsymbol{y}$. Note that due to $\Pi_{x}\Pi_{y}=\delta_{x,y}\Pi_{x}$, it is always true that $\mathfrak{D}(\boldsymbol{x};\boldsymbol{y})\sim\delta_{x_{n},y_{n}}$, i.e, the DF is always “diagonal” with respect to the final points of the history. However, it is usually not true that the DF is also diagonal with respect to earlier times $t_{n-1},\cdots,t_{0}$ of the history. The special case where this holds for all times,
@@
is known as the decoherent histories condition (DHC). Then, only the diagonal elements of the DF survive, which equal the probability $\mathfrak{D}(\boldsymbol{x};\boldsymbol{x})$ to get measurement outcomes $\boldsymbol{x}$ according to Born’s rule. As a result, interference effects vanish and the coherent superposition $|\psi_{n}\rangle=\sum_{\boldsymbol{x}}|\psi(\boldsymbol{x})\rangle$ effectively behaves like the incoherent mixture $\rho(t_{n})=\sum_{\boldsymbol{x}}|\psi(\boldsymbol{x})\rangle\langle\psi(\boldsymbol{x})|$ (for a more detailed account highlighting the importance of the DHC see, e.g., Refs. [45, 1, 2, 6, 7, 3, 4, 46, 21, 47]).
@@
Figure 1: The expectation value $\langle A\rangle(t)$ of the coarse spin imbalance as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for the chaotic (i), interacting integrable (ii) and non-interacting integrable (iii) cases. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$.
@@
where $M^{2L-1}-M^{L}$ equals the number of non-trivial pairs $(\boldsymbol{x};\boldsymbol{y})$ (excluding those for which $\boldsymbol{x}=\boldsymbol{y}$ and $x_{n}\neq y_{n}$). Second, to capture the behaviour of statistical outliers we quantify the worst case scenario (the maximum coherence between histories) by
@@
Model.—As a paradigmatic and non-trivial quantum many-body system we consider a XXZ Heisenberg spin chain with Hamiltonian
@@
where $s^{\ell}_{x,y,z}=\sigma^{\ell}_{x,y,z}/2$ are spin operators at lattice sites $\ell$, $N$ is the length of the chain, and we assume periodic boundary conditions. Crucial for our purposes is that Eq. (8) contains three classes of systems for different parameter regimes: (i) for $\Delta_{1}\neq 0\neq\Delta_{2}$ (we choose $\Delta_{1}=1.5$, $\Delta_{2}=0.5$) the model is non-integrable (or chaotic) meaning that nearest-level-spacing follows a Wigner-Dyson distribution [48], and it satisfies the eigenstate thermalization hypothesis [49, 50]; (ii) for $\Delta_{1}\neq 0$ but $\Delta_{2}=0$ (we choose $\Delta_{1}=1.5$) it is an interacting integrable model, which is solvable by Bethe ansatz [51]; (iii) for $\Delta_{1}=\Delta_{2}=0$ the model is non-interacting integrable (or free) and can be mapped to a quadratic Hamiltonian (a set of free fermions) via Jordan-Wigner and Bogoliubov transformation [52]. Note that the fact that (ii) is integrable but can not be mapped to free fermions like (iii) is crucial: it qualitatively influences its transport behavior [53], operator complexity [54] and, as we reveal below, its decoherence.
@@
The emergence of classicality is investigated in Figs. 2 and 3 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$; the case $L=2$ has a universal typical decay owing to the Haar random nature of the initial state [20] as exemplified in the supplemental material (SM). The plots are obtained for constant time intervals $t_{k}-t_{k-1}=T$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. 2 (identical to the dotted line in Fig. 1) and an equilibrium time scale $T_{\text{eq}}$ in Fig. 3 (identical to the dash-dotted line in Fig. 1). More precisely, $T_{\text{neq}}$ is defined as the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value [55] and the equilibrium time scale is defined as $T_{\text{eq}}=20T_{\text{neq}}$. While in the free model (iii) there is no clearly visible equilibration in Fig. 1, we use the same convention for $T_{\text{eq}}$ for better comparison. Finally, each data point in Figs. 2 and 3 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{30-N}$ different realizations of $|\psi_{0}\rangle$. In the SM we show that the variance of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ decays as $D^{-1}$ for all three cases (i), (ii) and (iii) (likely as a consequence of dynamical typicality [56, 57]), which justifies our focus on the average behaviour.
@@
We then observe in Figs. 2 and 3 the following (an explanation follows later). First, for the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D$ scales exponentially with the number of spins $N$) with $\alpha\approx 0.35$ ($\alpha\approx 0.25$) for $\overline{\epsilon}$ ($\epsilon_{\text{max}}$) at the nonequilibrium time scale and with $\alpha\approx 0.5$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ at the equilibrium time scale. This indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the interacting integrable case (ii) we find three major differences compared to (i). First, we can safely conclude that the exponent $\alpha$ is smaller in all cases. Second, for $T_{\text{eq}}$ $\alpha$ is smaller than for $T_{\text{neq}}$, conversely to case (i). Indeed, for $T_{\text{eq}}$ (Fig. 3) we even find $\alpha\approx 0$, indicating a possible sub-exponential or power-law suppression (with respect to $N$) of coherences. Third, the exponent $\alpha$ for $\epsilon_{\text{max}}$ is roughly half the magnitude than $\alpha$ for $\overline{\epsilon}$ at $T_{\text{neq}}$, indicating stronger fluctuations in the DF among different pairs $(\boldsymbol{x};\boldsymbol{y})$ of histories. Finally, the free case (iii) is characterized for $T_{\text{neq}}$ by much larger coherences and for $T_{\text{eq}}$ by strong fluctuations among different $N$, making it unjustified to even speak of any scaling law $D^{-\alpha}$.
Figure 4: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ for history length $L=5$ for chaotic (a,d), interacting integrable (b,e) and non-interacting integrable (c,f) cases for time steps $T=T_{\text{neq}},2T_{\text{neq}},3T_{\text{neq}},T_{\text{eq}}$ (from light to dark blue). The dashed line indicate the scaling $D^{-1/2}$. The system size is $N=18,20,\dots,28$.
A particularly intriguing observation is the distinctive behaviour of decoherence as a function of $T$, which is further studied in Fig. 4. Whereas in case (i) $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ consistently decrease and the exponent grows to $\alpha\approx 0.5$ for increasing $T$, the opposite happens for case (ii): decoherence becomes consistently weaker for larger $T$ and no exponential scaling is visible anymore for $\epsilon_{\text{max}}$ at $T_{\text{eq}}$. Finally, for case (iii) no clear pattern is visible and the large fluctuations make it difficult to speak of any scaling at all.
@@
Explanation.—We start by noting that there is some consensus about the qualitative origin of decoherence (whether in the histories or EID framework), namely that coarse and slow (or “quasi-conserved”) observables of many-body systems behave classical. However, other qualitative questions (e.g., is non-integrability essential or not?) have not been addressed in the past, and useful estimates of the DF are hard to obtain in general [20, 21]. Nevetherless, it seems that al least a transparent explanation for the quantitative behaviour of the DF for chaotic systems can be given.
@@
To this end, recall that the overlap $\langle\phi|\chi\rangle$ between two Haar random vectors $|\phi\rangle$ and $|\chi\rangle$ scales like $1/\sqrt{D}$ and that the subspace dimensions $V_{x}$ are proportional to $D$ for many relevant coarse-grainings. Thus, for equilibrium time scales $T_{\text{eq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ in a chaotic system behave like randomly drawn typical states because they had time to explore the available Hilbert space in an unbiased way owing to the absence of conserved quantities (besides energy). Now, for nonequilibrium time scales $T_{\text{neq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ will not look fully randomized as they contain further information compared to the equilibrium case. Therefore, the exponent $\alpha=\alpha(T)$ is expected to grow with $T$ until it reaches $\alpha(T)\approx 0.5$ for $T\gtrsim T_{\text{eq}}$. Which values it takes precisely is, however, determined by $H$, $\{\Pi_{x}\}$ and $|\psi_{0}\rangle$ in a complicated way.
@@
Unfortunatey, it seems much more complicated to explain the behaviour for cases (ii) and (iii). Certainly, owing to the extensive number of conserved quantities, the states $|\psi(\boldsymbol{x})\rangle$ can not explore the available Hilbert space in an unbiased fashion, which causes deviations from the behaviour of typical states. Yet, why the exponent $\alpha(T)$ becomes even smaller for larger $T$ remains a mystery. A speculative guess could be that at short time scales the dynamics is governed by time-dependent perturbation theory (Fermi’s golden rule), which is more determined by the rough structure of the spectrum of the Hamiltonian. At long times, instead, the fine structure of the Hamiltonian (its conserved quantities) become increasingly important for the dynamical behaviour.
@@
Conclusion.—We numerically extracted finite size scaling laws for the DF of a realistic quantum many-body system in an approximation-free way and we revealed decisive differences depending on the nature of the system. The chaotic case (i) showed a strong and robust emergence of decoherence in contrast to the interacting integrable case (ii) with a quantitatively much weaker form of decoherence. Note that this difference could not have been guessed from the quantitatively almost identical single-time behaviour shown in Fig. 1. A qualitative even weaker form of decoherence was observed for free models (iii), making it even hard to speak of any definite signature of decoherence, even though our finite size calculations do not directly invalidate conclusions obtained from free models in the thermodynamic limit [5, 6, 7, 8, 10, 11, 12]. In particular, for another model (energy exchanges in an Ising chain, studied in the SM) we found a clear signature of decoherence also for the free case (while still being much weaker compared to the chaotic case).
@@
Acknowledgement.—JW is financially supported by the Deutsche Forschungsgemeinschaft (DFG), under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190. PS is financially supported by “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/PR21/11840014) as well as the MICINN with funding from European Union NextGenerationEU (PRTR-C17.I1), by the Generalitat de Catalunya (project 2017-SGR-1127), the European Commission QuantERA grant ExTRaQT (Spanish MICIN project PCI2022-132965), and the Spanish MINECO (project PID2019-107609GB-I00) with the support of FEDER funds. Additionally, we greatly acknowledge computing time on the HPC3 at the University of Osnabrück, granted by the DFG, under Grant No. 456666331.
@@
We here report on further numerical results that strengthen our main message. We start by supplementing additional results for the XXZ Heisenberg chain studied in the main text (subsection S1), continued by comparing the chaotic and free case for energy exchanges in an Ising chain (subsection S2), and finally we look at the level statistics to demonstrate that our models are chaotic (subsection S3).
.1 S1. Additional numerical results in XXZ model
We begin by showing plots similar to Figs. 2 and 3 in the main text, but for time-interval $T=2T_{\text{neq}}$ and $T=3T_{\text{neq}}$. Since $T_{\text{neq}}<T<T_{\text{eq}}$, we expect that Figs. S1 and S2 interpolates between the conclusions we obtained from Figs. 2 and 3, which is indeed the case. We find a stronger exponential suppression of decoherence for case (i) and a weaker exponential suppression for case (ii). Interestingly, the magnitude of decoherence for case (iii) is similar to case (ii) in contrast to Fig. 2, but it starts to fluctuate more, making a fit to a scaling law $D^{-\alpha}$ unreasonable as in Fig. 3.
Figure S1: Identical to Figs. 2 and 3 in the main text except for time steps $T=2T_{\text{neq}}$.
Figure S2: Identical to Figs. 2 and 3 in the main text except for time steps $T=3T_{\text{neq}}$.
In Fig. S3 we display $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ for the shortest possible non-trivial history with length $L=2$. As claimed in the main mansuscript, we observe consistently a $D^{-0.5}$ scaling in all three cases (i), (ii) and (iii). This effect is due to the Haar random choice of the initial state and the fact that the Haar measure is invariant under unitary transformations. Thus, the first unitary time evolution $U_{1,0}$ is barely able to change the decoherence properties of the system.
@@
As a start, we plot the expectation value $\langle B\rangle(t)$ of the coarse energy imbalance as a function of time in Fig. S8. This is done for two different non-equilibrium initial states $|\psi_{0}^{\pm}\rangle$, where $|\psi_{0}^{\pm}\rangle$ is a Haar random state restricted to the subspace ±${\cal{H}}^{B}_{\pm}$. In the chaotic case (i) the system relaxes to its thermal equilibrium value $\langle B\rangle_{\text{eq}}=0$ with an equilibration time scale $\propto N^{2}$. In contrast, in the free case (ii) $\langle B\rangle(t)$ decays on a time scale $\propto N$ and fluctuates around $\langle B\rangle_{\text{eq}}$ without any clearly visible equilibration (up to the time that we considered). Note, however, that—compared to the free case of the Heisenberg chain considered in the main text [cf. Fig. 1]—equilibration seems to work much better for the free case of the Ising model.

### v2

@@ Abstract
Introduction.—The decoherence functional (DF)—quantifying the (de)coherence between different paths or histories in an isolated quantum system (precisely defined below)—plays a crucial role to explain the emergence of classicality [1, 2, 3, 4]. Fundamentally, it provides the condition for the existence of records of past events [5, 6, 7, 8, 9, 10], which may be used to identify branches in the universal wave function [1, 11], and it implies macrorealism (Leggett-Garg inequalities) [12, 13]. Practically, the DF determines whether a complex quantum process can be simulated by a simpler classical stochastic process [14, 15, 16, 17, 18, 19, 20, 21].
Historically, since the DF is a complicated multi-time correlation function, research has been restricted to evaluating it explicitly for the case of quantum Brownian motion only [22, 23, 6, 7, 24, 25, 26]. This non-interacting integrable model has been solved for a bath prepared in a canonical ensemble, a state that is highly mixed and contains (in the conventionally considered continuum limit) an infinite amount of classical noise. This leaves it unclear whether the microscopic origin of the observed decoherence is an intrinsic feature of the dynamics or an artifact of the classical initial state. Also other indirect arguments for emergent decoherent histories were based on linear oscillator chains [27]. In addition, the thermodynamic limit makes the question how exactly decoherence emerges as a function of system size inaccessible.
@@
Our results also illuminate the debated relation to environmentally induced decoherence (EID) [37, 38, 39], which is mathematically not equivalent to decoherent histories (for detailed studies see Refs. [40, 5, 41, 42, 9, 20]). In our model, the relevant reduced density matrix exactly commutes with the relevant observable at all times, yet the histories are not exactly decoherent as one might naively expect. Since the block-diagonal form of the reduced density matrix is here caused by symmetry, this does not contradict the idea of EID, but it illustrates how subtle the relation is: it is a clear-cut example for emergent decoherent histories that are not caused by the entanglement between subsystems. Moreover, while (single-particle) chaos was found to be an obstacle for the quantum-to-classical transition, which can be cured by EID [43, 44, 45, 46, 47], this letter reveals that (many-body) chaos significantly enhances the emergence of decoherent histories, highlighting an intriguing dual role of chaos.
@@
Preliminaries.—We consider an isolated quantum system with Hamiltonian $H$, Hilbert space ${\cal H}$ with dimension $D=\dim{\cal{H}}$ and initial state $|\psi_{0}\rangle$. We divide the Hilbert space ${\cal{H}}=\bigoplus_{x}{\cal{H}}_{x}$ into orthogonal subspaces ${\cal{H}}_{x}$ corresponding to a complete set of orthogonal projectors $\{\Pi_{x}\}_{x=1}^{M}$ satisfying $\Pi_{x}\Pi_{y}=\delta_{x,y}\Pi_{x}$ and $\sum_{x=1}^{M}\Pi_{x}=I$ (with $I$ the identity). We call the set $\{\Pi_{x}\}$ a coarse-graining because the emergence of decoherence requires $\Pi_{x}$ to belong to a coarse observable $A=\sum_{x=1}^{M}a_{x}\Pi_{x}$ characterized by large subspace dimensions $V_{x}\equiv\dim{\cal{H}}_{x}\gg 1$.
@@
Here, $U_{k,i}=e^{-iH(t_{k}-t_{i})}$ is the unitary time evolution operator from time $t_{i}$ to $t_{k}$ ($\hbar\equiv 1$). For brevity, we denote a history as $\boldsymbol{x}=(x_{n},\dots,x_{1},x_{0})$ such that Eq. (1) becomes $|\psi_{n}\rangle=\sum_{\boldsymbol{x}}|\psi(\boldsymbol{x})\rangle$. Moreover, the length of a history is $L=n+1$.
@@
which quantifies the overlap, or interference, between different histories $\boldsymbol{x}$ and $\boldsymbol{y}$. Owing to $\Pi_{x}\Pi_{y}=\delta_{x,y}\Pi_{x}$, it is true that $\mathfrak{D}(\boldsymbol{x};\boldsymbol{y})\sim\delta_{x_{n},y_{n}}$, i.e, the DF is always “diagonal” with respect to the final points of the history, but the DF is usually not diagonal with respect to earlier times $t_{n-1},\cdots,t_{0}$ of the history. The special case where
@@
is known as the decoherent histories condition (DHC). Then, only the diagonal elements of the DF survive, which equal the probability $\mathfrak{D}(\boldsymbol{x};\boldsymbol{x})$ to get measurement outcomes $\boldsymbol{x}$ according to Born’s rule.
@@
Figure 1: $\langle A\rangle(t)$ as a function of rescaled time for initial states $|\psi^{+}_{0}\rangle$ (solid line) and $|\psi^{-}_{0}\rangle$ (dashed line) for the chaotic (i), interacting integrable (ii) and non-interacting integrable (iii) cases. The dotted (dash-dotted) vertical line represents $T=T_{\text{neq}}$ ($T=T_{\text{eq}}$), where $T_{\text{neq}}$ indicates the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value and $T_{\text{eq}}=20T_{\text{neq}}$.
@@
where $M^{2L-1}-M^{L}$ equals the number of non-trivial pairs $(\boldsymbol{x};\boldsymbol{y})$ (excluding those for which $\boldsymbol{x}=\boldsymbol{y}$ and $x_{n}\neq y_{n}$). Second, statistical outliers and the worst case scenario (the maximum coherence between histories) are captured by
@@
Model.—As a paradigmatic quantum many-body system we consider a XXZ Heisenberg spin chain with Hamiltonian
@@
where $s^{\ell}_{x,y,z}=\sigma^{\ell}_{x,y,z}/2$ are spin operators at lattice sites $\ell$, $N$ is the length of the chain, and we assume periodic boundary conditions. Crucial for our purposes is that Eq. (8) contains three classes of systems for different parameter regimes: (i) for $\Delta_{1}\neq 0\neq\Delta_{2}$ (we choose $\Delta_{1}=1.5$, $\Delta_{2}=0.5$) the model is non-integrable (or chaotic) meaning that nearest-level-spacing follows a Wigner-Dyson distribution [60], and it satisfies the eigenstate thermalization hypothesis [61, 62]; (ii) for $\Delta_{1}\neq 0$ but $\Delta_{2}=0$ (we choose $\Delta_{1}=1.5$) it is an interacting integrable model, which is solvable by Bethe ansatz [63]; (iii) for $\Delta_{1}=\Delta_{2}=0$ the model is non-interacting integrable (or free) and can be mapped to a quadratic Hamiltonian (a set of free fermions) via Jordan-Wigner and Bogoliubov transformation [64]. The fact that (ii) is integrable but can not be mapped to free fermions like (iii) is crucial: it qualitatively influences its transport behavior [65], operator complexity [66] and, as we reveal below, its decoherence.
@@
Decoherence is investigated in Figs. 2 and 3 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$; the case $L=2$ has a universal typical decay owing to the Haar random nature of the initial state [19] as exemplified in the supplemental material (SM). The plots are obtained for constant time intervals $t_{k}-t_{k-1}=T$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. 2 (identical to the dotted line in Fig. 1) and an equilibrium time scale $T_{\text{eq}}$ in Fig. 3 (identical to the dash-dotted line in Fig. 1). More precisely, $T_{\text{neq}}$ is defined as the time at which $\langle A\rangle(t)$ decays to $e^{-1}$ of its initial value [67] and the equilibrium time scale is defined as $T_{\text{eq}}=20T_{\text{neq}}$. While in the free model (iii) there is no clearly visible equilibration in Fig. 1, we use the same convention for $T_{\text{eq}}$ for better comparison. Finally, each data point in Figs. 2 and 3 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{30-N}$ different realizations of $|\psi_{0}\rangle$. In the SM we show that the variance of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ decays as $D^{-1}$ for all three cases (i), (ii) and (iii) (likely as a consequence of dynamical typicality [68, 69]), which justifies our focus on the average behaviour.
@@
We then observe in Figs. 2 and 3 the following (an explanation follows later). First, for the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D$ scales exponentially with the number of spins $N$) with $\alpha\approx 0.35$ ($\alpha\approx 0.25$) for $\overline{\epsilon}$ ($\epsilon_{\text{max}}$) at the nonequilibrium time scale and with $\alpha\approx 0.5$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ at the equilibrium time scale. This indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the interacting integrable case (ii) we find three major differences compared to (i). First, the exponent $\alpha$ is notably smaller in all cases. Second, for $T_{\text{eq}}$ $\alpha$ is smaller than for $T_{\text{neq}}$, conversely to case (i). Indeed, for $T_{\text{eq}}$ (Fig. 3) we even find $\alpha\approx 0$, indicating a possible sub-exponential or power-law suppression (with respect to $N$) of coherences. Third, the exponent $\alpha$ for $\epsilon_{\text{max}}$ is roughly half the magnitude than $\alpha$ for $\overline{\epsilon}$ at $T_{\text{neq}}$, indicating stronger fluctuations in the DF among different pairs $(\boldsymbol{x};\boldsymbol{y})$ of histories. Finally, the free case (iii) is characterized for $T_{\text{neq}}$ by much larger coherences and for $T_{\text{eq}}$ by strong fluctuations among different $N$, making it unjustified to even speak of any scaling law $D^{-\alpha}$.
Figure 4: Decoherence as a function of the time step $T=t_{n}-t_{n-1}$ for history length $L=5$. (a,b) $\overline{\epsilon},\epsilon_{\text{max}}$ for $N=24$ on a double logarithmic scale. (c,d) Scaling exponent $\alpha$ for $\overline{\epsilon},\epsilon_{\text{max}}$ obtained for system sizes $N=18,20,\ldots,28$ and as a mean over $2^{30-N}$ realizations of a Haar random $|\psi_{0}\rangle$. Red dashed line in (c,d) indicates $\alpha=0.5$.
A particularly intriguing observation is the distinctive behaviour of decoherence as a function of $T$, which is further studied in Fig. S2. We first observe that decoherence is forbidden for $T\rightarrow 0$ owing to the quantum Zeno effect. Still, decoherence sets in on a short nonequilibrium time scale for (i) and (ii) and later for (iii). However, what we find surprising is that decoherence becomes weaker again with increasing $T$ for (ii) and (iii) but not for (i). Moreover, the scaling exponent $\alpha$ clearly shows a non-trivial behaviour as a function of $T$.
@@
Explanation.—We note that there is some consensus about the qualitative origin of decoherence (whether in the histories or EID framework), namely that coarse and slow (or “quasi-conserved”) observables of many-body systems decohere. However, other qualitative questions (e.g., is non-integrability essential or not?) have not been addressed in the past, and useful estimates of the DF are hard to obtain [19, 20]. Nevertheless, at least some quantitative features of the DF for chaotic systems seem to have a transparent explanantion.
@@
To this end, recall that the overlap $\langle\phi|\chi\rangle$ between two Haar random vectors $|\phi\rangle$ and $|\chi\rangle$ scales like $1/\sqrt{D}$ and that the subspace dimensions $V_{x}$ are proportional to $D$ for many relevant coarse-grainings. Thus, for equilibrium time scales $T_{\text{eq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ in a chaotic system behave like randomly drawn typical states because they had time to explore the available Hilbert space in an unbiased way owing to the absence of conserved quantities (besides energy). Instead, for nonequilibrium time scales $T_{\text{neq}}$ the history states $|\psi(\boldsymbol{x})\rangle$ will not look fully randomized as they contain further information compared to the equilibrium case. Therefore, one expects the exponent $\alpha=\alpha(T)$ to obey $\alpha(T)<0.5$ for $T<T_{\text{eq}}$, but Fig. S2 reveals exceptions for short time windows. This indicates that the precise time dependence of $\alpha(T)$ is determined by $H$, $\{\Pi_{x}\}$ and $|\psi_{0}\rangle$ in a complicated way.
@@
Unfortunatey, it is even more complicated to explain the behaviour for cases (ii) and (iii). Certainly, owing to the extensive number of conserved quantities, the states $|\psi(\boldsymbol{x})\rangle$ can not explore the available Hilbert space in an unbiased fashion, which causes deviations from the behaviour of typical states. Yet, why the exponent $\alpha(T)$ becomes even smaller for larger $T$ remains a mystery. A speculative guess could be that at short time scales the dynamics is governed by time-dependent perturbation theory (Fermi’s golden rule), which is more determined by the rough structure of the spectrum of the Hamiltonian. At long times, instead, the fine structure of the Hamiltonian (its conserved quantities) become increasingly important for the dynamical behaviour.
@@
Conclusion.—We numerically extracted finite size scaling laws for the DF of a realistic quantum many-body system in an approximation-free way and we revealed decisive differences depending on the nature of the system. The chaotic case (i) showed a strong and robust emergence of decoherence in contrast to the interacting integrable case (ii) with a quantitatively much weaker form of decoherence. Note that this difference could not have been guessed from the quantitatively almost identical single-time behaviour shown in Fig. 1. A qualitative even weaker form of decoherence was observed for free models (iii), making it even hard to speak of any definite signature of decoherence. Clearly, our finite size calculations do not directly invalidate conclusions obtained from free models in the thermodynamic limit [22, 23, 6, 7, 25, 26, 27]. In particular, for another model (energy exchanges in an Ising chain, studied in the SM) we found a clear signature of decoherence also for the free case (while still being much weaker compared to the chaotic case).
@@
The notable difference in the value $\alpha$ at equilibrium time scales (and its time dependence) between chaotic and integrable systems suggests that $\alpha$ can be used as an indicator of quantum chaos. A main advantage of it is the applicability to system sizes beyond the reach of exact diagonalization, e.g., through real-time propagation methods such as the Chebyshev polynomial algorithm used here.
@@
Acknowledgement.—JW is financially supported by the Deutsche Forschungsgemeinschaft (DFG), under Grant No. 531128043, as well as under Grant No. 397107022, No. 397067869, and No. 397082825 within the DFG Research Unit FOR 2692, under Grant No. 355031190. PS is financially supported by “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/PR21/11840014), the Ramón y Cajal program RYC2022-035908-I, the MICINN with funding from European Union NextGenerationEU (PRTR-C17.I1), by the Generalitat de Catalunya (project 2017-SGR-1127), the European Commission QuantERA grant ExTRaQT (Spanish MICIN project PCI2022-132965), and the Spanish MINECO (project PID2019-107609GB-I00) with the support of FEDER funds. Additionally, we greatly acknowledge computing time on the HPC3 at the University of Osnabrück, granted by the DFG, under Grant No. 456666331.
@@
S1. Additional numerical results in XXZ model
We begin in Fig. S1 by displaying $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ for the shortest possible non-trivial history with length $L=2$. As claimed in the main mansuscript, we observe consistently a $D^{-0.5}$ scaling in all three cases (i), (ii) and (iii). This effect is due to the Haar random choice of the initial state and the fact that the Haar measure is invariant under unitary transformations. Thus, the first unitary time evolution $U_{1,0}$ is barely able to change the decoherence properties of the system.
@@
Figure S2: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus Hilbert space dimension $D$ for history length $L=5$ for chaotic (a,d), interacting integrable (b,e) and non-interacting integrable (c,f) cases for time steps $T=T_{\text{neq}},2T_{\text{neq}},3T_{\text{neq}},T_{\text{eq}}$ (from light to dark blue). The dashed line indicate the scaling $D^{-1/2}$. The system size is $N=18,20,\dots,28$.
In addition to Fig. 4 in the main text, we further study the distinctive behaviour of decoherence as a function of $T$ in Fig. S2. Whereas in case (i) $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ consistently decrease and the exponent grows to $\alpha\approx 0.5$ for increasing $T$, the opposite happens for case (ii): decoherence becomes consistently weaker for larger $T$ and no exponential scaling is visible anymore for $\epsilon_{\text{max}}$ at $T_{\text{eq}}$. Finally, for case (iii) no clear pattern is visible and the large fluctuations make it difficult to speak of any scaling at all.
@@
As a start, we plot the expectation value $\langle B\rangle(t)$ of the coarse energy imbalance as a function of time in Fig. S7. This is done for two different non-equilibrium initial states $|\psi_{0}^{\pm}\rangle$, where $|\psi_{0}^{\pm}\rangle$ is a Haar random state restricted to the subspace ±${\cal{H}}^{B}_{\pm}$. In the chaotic case (i) the system relaxes to its thermal equilibrium value $\langle B\rangle_{\text{eq}}=0$ with an equilibration time scale $\propto N^{2}$. In contrast, in the free case (ii) $\langle B\rangle(t)$ decays on a time scale $\propto N$ and fluctuations around $\langle B\rangle_{\text{eq}}$ remain visible for all times that we considered. Note, however, that—compared to the free case of the Heisenberg chain considered in the main text [cf. Fig. 1]—equilibration seems to work much better for the free case of the Ising model.